Annales Academiæ Scientiarum Fennicæ Mathematica Volumen 25, 2000, 501–522

CALORIC MEASURE ON DOMAINS BOUNDED BY WEIERSTRASS-TYPE GRAPHS

Thierry Bousch and Yanick Heurteaux

Université Paris-Sud, Département de Mathématique Bât. 425-430, F-91405 Orsay Cedex, France Thierry.Bousch@math.u-psud.fr, Yanick.Heurteaux@math.u-psud.fr

Abstract. Let ω be the caloric measure on a domain bounded by a Weierstrass-type graph. We prove that ω is a quasi-Bernoulli measure. Therefore, the multifractal formalism is available for such a measure. We can then give necessary and sufficient conditions which ensure that ω is supported by a set of small dimension and prove that this is the generic case. These results improve and generalize an earlier work due to R. Kaufman and J.M. Wu.

Let f: R → R be a real function of class C^{1/2} (i.e., Hölder of exponent 1/2) and Ω = {(x, t) ∈ R^{2} : x > f(t)}. For every point M_{0} = (x_{0}, t_{0}) ∈ Ω, let ω^{M_{0}} be the harmonic measure at M_{0} with respect to the heat operator 𝒞 = ∂/∂t−∂^{2}/2∂x^{2} . This measure is called the caloric measure at M_{0} . It is supported by the set of points (x, t) ∈ ∂Ω such that t ≤ t_{0} and may also be defined by

ω^{M_{0}}(E) = P_{x_{0}}(T < +∞ and (B_{T} , t_{0} − T) ∈ E),

where B_{t} is a standard Brownian motion and

T = inf({t > 0 : (B_{t}, t_{0} − t) ∈ R^{2} \ Ω}).

We are interested in the geometric properties of the measure ω^{M_{0}} . The domain Ω can be seen as a Lipschitz domain related to the parabolic distance in R^{2}

δ((x, t), (y, s)) = sup(|x − y|, |t − s|^{1/2}).

Because of the homogeneity properties of the heat operator, this distance is adapted to our situation. It is then natural to ask whether the analogue of Dahlberg’s theorem ([Da]) holds for the heat equation in Ω. Of course, the graph of a function of class C^{1/2} is in general not rectifiable. Nevertheless, following Taylor and Watson ([TW]), we can introduce the Hausdorff measures related to the metric space (R^{2}, δ). These are defined for α > 0 by

Λ_{α}(E) = lim_{ε→0}(inf(∑_{i} r_{i}^{α}, E ⊂ ⋃_{i} D(a_{i}, r_{i}) and r_{i} ≤ ε)),

1991 Mathematics Subject Classification: Primary 31B25, 35K05; Secondary 28A80, 28A78.

where D(a_{i}, r_{i}) are balls for the distance δ. The related dimension is called the parabolic dimension. In this context, the parabolic dimension of the graph ∂Ω is equal to 2 and its Λ_{2} -measure is locally finite. Then, there is an analogy between the measure Λ_{2} in graphs of C^{1/2} -functions and the length measure in Lipschitz graphs.

In [KW], Robert Kaufman and Jang-Mei Wu proved that the equivalent of Dahlberg’s theorem is false for the heat equation and constructed a function f of class C^{1/2} such that the caloric measure is supported by a set of parabolic dimension strictly less than 2. In other words, if ω̃^{M_{0}} denotes the image of the caloric measure ω^{M_{0}} under the projection

(1)  Π: (f(t), t) ∈ ∂Ω ⟼ t ∈ R,

the measure ω̃^{M_{0}} is supported by a set of Hausdorff dimension strictly less than 1.

On the other hand, for slightly more regular functions f, the measure ω̃^{M_{0}} becomes equivalent to the Lebesgue measure dt. For example, in [LS], Lewis and Silver proved that this is the case when the modulus of continuity φ of the function f satisfies the Dini condition ∫_{0}^{1} φ^{2}(t)/t^{2} dt < +∞. Moreover, when f is of class C^{(1+ε)/2} , we can prove that the relative densities between the measures ω̃^{M_{0}} and dt are locally bounded functions (see [H2, pp. 642–643]).

In this work, we want to give a more precise analysis of the caloric measure on domains bounded by Weierstrass-type graphs. Let g be a Lipschitz function, periodic with period 1, and l be an integer greater than 1. We are interested in the caloric measure on the domain Ω bounded by the graph of the function

(2)  f(t) = ∑_{k=0}^{+∞} l^{−k/2}g(l^{k}t).

Such a curve ∂Ω will be called a Weierstrass-type curve (see [F]) and the domain Ω a Weierstrass domain. It is well known that when the function g is Lipschitz with Lipschitz constant K̃, the function f is of class C^{1/2} with Hölder constant K = 2K̃/(1 − l^{−1/2}). In other words, the map g ↦ f is continous from the space of 1-periodic Lipschitz functions to the space of 1-periodic functions of class C^{1/2} when we endow these spaces with their natural norms.

The first result of this work (Theorem 3.1) states that for a Weierstrass domain Ω and for suitable points M_{0} , the measure ω̃^{M_{0}} is a quasi-Bernoulli measure in [0, 1) (see (16)). To prove this theorem, we need a preliminary result which is developed in Section 2. When ∂Ω_{2} is a Lipschitz perturbation of ∂Ω_{1} , we establish in Theorem 2.1 that the associated caloric measures (more precisely their projections on the real axis) are strongly equivalent.

Quasi-Bernoulli measures have been extensively studied in the literature ([C], [MV], [BMP], [H5], [BH]) and we can use known results to describe the geometric properties of the measure ω̃^{M_{0}} . Let ℱ_{n} be the family of l-adic intervals of [0, 1) of the n^{th} generation and denote by I_{n}(t) the unique interval of ℱ_{n} that contains t. It is well known (see for example [C] or [H5]) that there exists a real d ≤ 1 such that for almost every t,

lim_{n→+∞} \frac{log(ω̃^{M_{0}}(I_{n}(t)))}{−n log l} = d.

In particular, the measure ω̃^{M_{0}} is supported by a set of dimension d and every set of Hausdorff dimension < d is negligible (in this case, we say that the measure is unidimensional and we call d = dim (ω̃^{M_{0}}) the dimension of the measure). In fact, the number d can be calculated in the following way. Let λ = ω̃^{M_{0}}/ω̃^{M_{0}}([0, 1)) be the normalized measure and define

(3)  τ_{n}(x) = \frac{1}{n log l} log( ∑_{I∈ℱ_{n}} λ(I)^{x})  and  τ(x) = lim_{n→+∞} τ_{n}(x)

(the limit exists in such a situation). It is proved in [H5] that the function τ is differentiable and satisfies

(4)  dim(ω̃^{M_{0}}) = −τ′(1) = lim_{n→+∞} \frac{−1}{n log l} ∑_{I∈ℱ_{n}} λ(I) log(λ(I)).

Then we can conclude (see Corollary 3.2) that either the measure ω̃^{M_{0}} is equivalent to the Lebesgue measure on its support with locally bounded densities, or it has dimension d < 1.

Let us also remember that the multifractal formalism is available for such a measure. Using the result of Brown–Michon–Peyriere ([BMP]) and the differentiability of the function τ ([H5]), we have

dim (E_{α}) = τ^{∗}(α) for all α ∈ (−τ′(+∞), −τ′(−∞)),

where τ^{∗} is the Legendre transform of the convex function τ and

E_{α} = \{t ∈ [0, 1) : lim_{r→0} \frac{log(ω̃^{M_{0}}([t − r, t + r]))}{log r} = α\}.

In particular, in the case where dim (ω̃^{M_{0}}) < 1, we obtain that the measure ω̃^{M_{0}} is multifractal in the sense that the set E_{α} has a positive dimension for infinitely many values of α.

In Section 4, we characterize those functions g for which the dimension of the measure ω̃^{M_{0}} is strictly less than 1. Of course this is not the case for every Lipschitz function g. If there exists a Lipschitz function γ such that g(t) = γ(t) − l^{−1/2}γ(lt), then f = γ is Lipschitz and the measure ω̃^{M_{0}} is equivalent to the Lebesgue measure. In fact, this is exceptional. A careful study of the functional equation f(t) = g(t) + f(lt)/\sqrt{l} allows us to prove that dim (ω̃^{M_{0}}) = 1 if and only if f is Lipschitz. In particular, the set of Lipschitz functions g such that dim (ω̃^{M_{0}}) < 1 is an open dense subset of the space of Lipschitz functions. Moreover, in Proposition 4.7, we introduce an explicit criterion to test whether the function f is Lipschitz.

It is also interesting to give a lower bound for the dimension of the measure ω̃^{M_{0}} . It seems difficult to obtain an optimal lower bound. Using Taylor–Watson results ([TW]), we know that, when f is of class C^{1/2} , every set of dimension strictly less than 1/2 is negligible for ω̃^{M_{0}} (its projection on the graph is polar for the heat equation). In fact, in Section 5, we improve this result and establish that there exists a constant ε = ε(K) such that every set of dimension strictly less than \frac{1}{2}(1 + ε) is negligible for the measure ω̃^{M_{0}} . A reasonable conjecture would be that for every ε > 0, it is possible to construct a Weierstrass like function f for which the dimension d of the measure ω̃^{M_{0}} is lower than \frac{1}{2}(1 + ε). However, we do not know how to prove this result.

1. The boundary Harnack principle and some consequences

In this section, we introduce some notation and recall the boundary Harnack principle for the heat equation. We also state two important consequences of this principle. All of these results can be found in [H1] or in [H2] where they are proved for general parabolic operators. The domains we are working with are not necessarily of Weierstrass-type. Let f: R → R be a bounded function such that for some K > 0

(5)  |f(t) − f(t′)| ≤ K|t − t′|^{1/2}  for all (t, t′) ∈ R^{2}.

Put Ω = {(x, t) ∈ R^{2} : x > f(t)}. For every Q = (y, s) ∈ ∂Ω and r > 0 define

(6)  \{ \begin{matrix}T (Q, r) = \{(x, t) ∈ Ω : |t − s| < r, |x − y| < 10K\sqrt{r} \}, \\ Δ(Q, r) = \{(x, t) ∈ R^{2} : |t − s| < r, x = f(t)\}. \end{matrix}

The following comparison theorem is due to J.T. Kemper ([K]). A generalization is given for parabolic operators in [H1] and [H2].

Theorem 1.1. There exists a strictly positive constant C = C(K) such that for every non-negative caloric function u in Ω (i.e., satisfying 𝒞 u = 0) which converges to 0 on ∂Ω \ Δ(Q, \frac{1}{2}r) and which is dominated by a potential in a neighbourhood of infinity, we have

(7)  u(P) ≤ Cu(M_{r})  for all P ∈ Ω \ T (Q, r),

where M_{r} = Q + (10K\sqrt{r} , r).

Remarks. 1. Since the domain Ω is unbounded, we need a domination hypothesis. This hypothesis allows us to use the maximum principle. It is automatically satisfied when the function u can be written u(P) = ∫ ψ(M) dω^{P} (M) with ψ a bounded function supported by Δ(Q, \frac{1}{2}r).

2. The conclusions of Theorem 1.1 remain true when u is caloric in Ω \ T (Q, \frac{1}{2}r).

The two following results describe comparison principles near the boundary between two non-negative solutions of the heat equation and can be found in [H1] and [H2].

Theorem 1.2 (Weak boundary Harnack principle). There exists a strictly positive constant C = C(K) such that if u and v are two non-negative caloric functions in T (Q, 4r) which are equal to 0 at every point of Δ(Q, 2r), then

(8)  \frac{u(P)}{u(P_{r})} ≤ C \frac{v(P)}{v(P_{r}^{∗})}  for all P ∈ T (Q, r),

where P_{r} = Q + (10K\sqrt{r} , 2r) and P_{r}^{∗} = Q + (10K\sqrt{r} , −2r).

Theorem 1.3 (Strong boundary Harnack principle). There exists a strictly positive constant C = C(K) such that if u and v are two non-negative caloric functions in Ω \ T (Q, r) which are equal to 0 at every point of ∂Ω \ Δ(Q, r) and which are dominated by a potential in a neighbourhood of infinity, then

(9)  \frac{u(P_{s})}{u(P_{r})} ≤ C \frac{v(P_{s})}{v(P_{r})}  for all s > r.

Remark. In [H1] and [H2], inequalities (8) and (9) are established for a large class of parabolic operators and require respectively r ≤ 1 and s ≤ 1. Because of the homogeneity of the heat operator 𝒞 , these assumptions are not needed in the present situation. In [H2], inequality (9) is established for bounded domains. The general case can be found in [H1] and does not need additional ideas. Let us also recall that an inequality similar to (9) is established in [Ny] for bounded domains and for slightly more general operators, the constant C depending on diam (Ω).

An important consequence of the boundary Harnack principle is the doubling property satisfied by the caloric measure. This principle was observed by many authors (see for example [W], [FGS], [H1], [H4] or [Ny]). Let us remember what it means.

Theorem 1.4 (Doubling property for the caloric measure). Let Q_{0} ∈ ∂Ω, r_{0} > 0 and put M_{0} = Q_{0}+(10K\sqrt{r_{0}} , 2r_{0}). There exists a strictly positive constant C = C(K) such that if Δ(Q, 2r) ⊂ Δ(Q_{0}, \frac{1}{2}r_{0}), then,

(10)  ω^{M_{0}}(Δ(Q, 2r)) ≤ Cω^{M_{0}}(Δ(Q, r)).

We will end this section with another important consequence of the boundary Harnack principle which gives comparisons between the caloric measure and the Green function on Ω. These estimates were proved by Wu in [W] for the heat equation and extended to parabolic operators in [H2].

Theorem 1.5. Let Q_{0} ∈ ∂Ω, r_{0} > 0 and put M_{0} = Q_{0} + (10K\sqrt{r_{0}} , 2r_{0}). Denote by G(A, ·) the Green function of Ω related to the heat equation with singularity at A. There exists a strictly positive constant C = C(K) such that if Δ(Q, s) ⊂ Δ(Q_{0}, \frac{1}{2}r_{0}), then

(11)  \frac{1}{C} ω^{M_{0}}(Δ(Q, s)) ≤ \sqrt{s} G(Q_{s}, M_{0}) ≤ Cω^{M_{0}}(Δ(Q, s)),

where Q_{s} = Q + (\sqrt{s} , 0).

Notation. The important fact in (7), (8), (9), (10) and (11) is that the constants C are independent of r and Q. We write A ∼ < B when there exists such a constant C for which the quantities A and B satisfy A ≤ CB. The notation A ≈ B means that the relations A ∼ < B and B ∼ < A are true.

2. Behavior of the caloric measure under perturbations of ∂Ω

In this section, we establish a preliminary result which will be the key of the proof of Theorem 3.1. Let f_{1} and f_{2} be two bounded functions satisfying (5), a time t_{0} ∈ R and a real r > 0. For i = 1, 2, denote by Q_{i}(t) = (f_{i}(t), t), M_{i} = Q_{i}(t_{0}) + (20K\sqrt{r} , 8r) and ω_{i}^{M_{i}} the caloric measure related to the domain Ω_{i} = {x > f_{i}(t)} and to the point M_{i}. We have the following result:

Theorem 2.1. Let h = f_{2} − f_{1} and suppose that for some K′ > 0

(12)  |h(t) − h(t′)| ≤ \frac{K′}{\sqrt{r}} |t − t′|  for all t, t′ ∈ [t_{0} − r, t_{0} + r].

In other words, suppose that f_{2} is a Lipschitz perturbation of f_{1} in [t_{0} −r, t_{0} +r]. Let ω̃_{i} be the image of the measure ω_{i}^{M_{i}} under the projection (f_{i}(t), t) ↦ t. There exists a strictly positive constant C = C(K, K′) such that for every Borel set E ⊂ [t_{0} − r, t_{0} + r], we have

(13)  \frac{1}{C} ω̃_{2}(E) ≤ ω̃_{1}(E) ≤ Cω̃_{2}(E).

Proof. Let us begin with some reductions of the problem. Because of the regularity properties of the measures ω̃_{i}, we may assume that E is an open set and even an interval E = [s_{0}−s, s_{0}+s) of small length. Using the translation invariance of the heat equation, we may also assume that s_{0} = 0 and f_{1}(s_{0}) = f_{2}(s_{0}) = 0. Let

U_{i} = Ω_{i} ∩ (\{(x, t) ∈ R^{2} : |x| < 10K\sqrt{r} and |t| < r\})

and

𝒞^{∗} = −\frac{∂}{∂t} − \frac{∂^{2}}{2∂x^{2}}

the adjoint heat operator. Denote by φ_{i}^{∗} the 𝒞^{∗} -harmonic measure of ∂U_{i} \ ∂Ω_{i} in U_{i}. If Δ_{i} is the projection of E on ∂Ω_{i} and if G_{i}(A, ·) is the Green function of Ω_{i} with singularity at A, we deduce from Theorem 1.5 and from the weak boundary Harnack principle related to the adjoint heat operator that for small values of s,

\begin{aligned}ω_{i}^{P_{r}}(Δ_{i}) & ≈ \sqrt{s} G_{i}((\sqrt{s} , 0), P_{r}) \\ & ∼ < \sqrt{s} \frac{φ_{i}^{∗}((\sqrt{s} , 0))}{φ_{i}^{∗}(P_{r/4})} G_{i}(P_{r/4}^{∗}, P_{r}) ≈ \sqrt{\frac{s}{r}} φ_{i}^{∗}((\sqrt{s} , 0)) \end{aligned}

where P_{r} = (10K\sqrt{r} , 2r) and P_{r}^{∗} = (10K\sqrt{r} , −2r). Conversely, we have

\sqrt{s} G_{i}((\sqrt{s} , 0), P_{r}) ∼ > \sqrt{s} \frac{φ_{i}^{∗}((\sqrt{s} , 0))}{φ_{i}^{∗}(P_{r/4}^{∗})} G_{i}(P_{r/4}, P_{r}) ≈ \sqrt{\frac{s}{r}} φ_{i}^{∗}((\sqrt{s} , 0))

and we can conclude that

ω_{i}^{P_{r}}(Δ_{i}) ≈ \sqrt{\frac{s}{r}} φ_{i}^{∗}((\sqrt{s}, 0)).

On the other hand, Harnack inequalities and Theorem 1.1 give respectively

\{ \begin{matrix}ω_{i}^{P_{r}}(Δ_{i}) ∼ < ω_{i}^{M_{i}}(Δ_{i}), \\ ω_{i}^{M_{i}}(Δ_{i}) ∼ < ω_{i}^{P_{r}}(Δ_{i}) \end{matrix}

and we obtain

(14)  ω̃_{i}(E) = ω_{i}^{M_{i}}(Δ_{i}) ≈ \sqrt{\frac{s}{r}} φ_{i}^{∗}((\sqrt{s} , 0)).

Finally, the proof of Theorem 2.1 will be complete if we establish that

(15)  φ_{1}^{∗}((\sqrt{s}, 0)) ≈ φ_{2}^{∗}((\sqrt{s} , 0))

for small values of s. This is an easy consequence of a result proved in [H2]. Let us introduce the map H

H: (x, t) ∈ R^{2} ⟼ (\sqrt{r} x, rt) ∈ R^{2}.

The homogeneity properties of the operator 𝒞^{∗} ensure that the function φ_{i}^{∗} ∘ H is the 𝒞^{∗} -harmonic measure of H^{−1}(∂U_{i} \ ∂Ω_{i}) in H^{−1}(U_{i}). This domain is delimited by the function f̃_{i}(t) = f_{i}(rt)/r^{1/2} and we can deduce from (12) that

|f̃_{1}(t) − f̃_{2}(t)| ≤ 2K′|t|.

Then, we can make use of [H2, Théorème 2.2] which is also true for the operator 𝒞^{∗} and obtain (15) for small values of s. This gives the conclusion of Theorem 2.1.

3. The quasi-Bernoulli property for the caloric measure

In this section, we choose a Lipschitz function g with Lipschitz constant K̃, periodic with period 1 and we define f using (2). As recalled in the introduction, f is C^{1/2} with Hölder constant K = 2K̃/(1 − l^{−1/2}). Denote by ℳ the set of finite words constructed with the alphabet {0, …, l − 1}. If a = ε_{1} ⋯ ε_{n} ∈ ℳ , define

t_{a} = ∑_{i=1}^{n} ε_{i}l^{−i},  I_{a} = [t_{a}, t_{a} + l^{−n}),  M_{a} = (f(t_{a}), t_{a}) + (20Kl^{−n/2}, 8l^{−n}),

and let |a| = n denote the length of a. Take the convention that if a = ∅, then t_{∅} = 0 and |∅| = 0. Finally, let us write ab the concatenation of the words a and b.

It is now possible to state the quasi-Bernoulli property for the caloric measure.

Theorem 3.1. Let ω^{M_{∅}} be the caloric measure in Ω at M_{∅} and ω be the image of ω^{M_{∅}} under the projection (f(t), t) ↦ t. The measure ω is a quasi-Bernoulli measure in [0, 1). In other words, there exists a strictly positive constant C = C(K̃), such that

(16)  \frac{1}{C} ω(I_{a})ω(I_{b}) ≤ ω(I_{ab}) ≤ Cω(I_{a})ω(I_{b})  for all a, b ∈ ℳ .

We can deduce the following corollary.

Corollary 3.2. There are only two possible cases, mutually exclusive.

(i) There exists a constant κ > 0 such that σ/κ ≤ ω ≤ κσ where σ is the Lebesgue measure on [0, 1). In particular dim (ω) = 1;

or (ii) there exists a real d < 1 such that ω is supported by a subset of [0, 1) with

dimension d and such that every set of dimension < d is negligible for ω; the measure ω is unidimensional with dimension

dim (ω) = d < 1.

Moreover, in this second case, ω is a multifractal measure. That is, dim (E_{α}) > 0 for infinitely many α ∈ R, where

(17)  E_{α} = \{t ∈ [0, 1) : lim_{r→0} \frac{log ω([t − r, t + r])}{log r} = α\}.

Remarks. 1. In the next section, we will prove that case (ii) is satisfied if and only if the function f is not Lipschitz; this is the generic situation.

2. We observe that self-similar curves for the caloric measure play the same role as self-similar Cantor sets for the harmonic measure. Carleson proved that the harmonic measure of a self-similar Cantor set is a quasi-Bernoulli measure ([C]). In 1986, Makarov and Volberg stated that it is supported by a set of dimension strictly less than the dimension of its support ([MV]). Let us also remember that Batakis generalized Makarov–Volberg’s result to a large class of Cantor sets ([Ba]).

3.1. Proof of Theorem 3.1. Using the invariance of the heat equation under translations, we can suppose that g(0) = 0. For every a ∈ ℳ , let Δ_{a} be the projection of I_{a} on the graph ∂Ω. Fix a ∈ ℳ and let M̃_{∅} = (f(t_{a}), t_{a})+(20K, 8). Using Harnack inequalities and Theorem 1.1, we first remark that

ω^{_{M}^{∅}}(Δ_{ab}) ≈ ω^{M_{∅}}(Δ_{ab}) for all b ∈ ℳ .

Then, by the strong boundary Harnack principle, we obtain

\frac{ω(I_{ab})}{ω(I_{a})} = \frac{ω^{M_{∅}}(Δ_{ab})}{ω^{M_{∅}}(Δ_{a})} ≈ \frac{ω^{_{M}^{∅}}(Δ_{ab})}{ω^{_{M}^{∅}}(Δ_{a})} ≈ \frac{ω^{M_{a}}(Δ_{ab})}{ω^{M_{a}}(Δ_{a})} ≈ ω^{M_{a}}(Δ_{ab}).

On the other hand, if n = |a|, we have

f(t) = ∑_{k=0}^{n−1} l^{−k/2}g(l^{k}t) + l^{−n/2}f(l^{n}t) = h(t) + l^{−n/2}f(l^{n}t),

where h is a Lipschitz function with Lipschitz constant

c ≈ l^{n/2} ≈ \frac{1}{\sqrt{σ(I_{a})}}.

Let us write with a hat all quantities related to the function f̂(t) = l^{−n/2}f(l^{n}t) and the domain Ω̂ = {(x, t) : x > f̂(t)}. Theorem 2.1 ensures that

ω^{M_{a}}(Δ_{ab}) ≈ ω̂^{_{M}^{a}}(Δ̂_{ab}).

If we observe that the domain Ω is the image of Ω̂ under the application

H: (x, t) ⟼ (l^{n/2}x, l^{n}t),

the invariance of the heat operator under such a transformation ensures that

ω̂^{_{M}^{a}}(Δ̂_{ab}) = ω^{H(_{M}_{a})}(H(Δ̂_{ab})).

Finally, observing that n_{0} = l^{n}t_{a} is an integer such that

H(M̂_{a}) = (0, n_{0}) + M_{∅}  and  H(Δ̂_{ab}) = (0, n_{0}) + Δ_{b},

and using the periodicity of the function f, we conclude that

ω^{H(_{M}_{a})}(H(Δ̂_{ab})) = ω^{M_{∅}}(Δ_{b}) = ω(I_{b}).

This completes the proof of Theorem 3.1.

3.2. Proof of Corollary 3.2. Corollary 3.2 is always true for quasi-Bernoulli measures. Let us begin with some notation. Let ℱ_{n} be the set of l-adic intervals of the n^{th} generation. If I = I_{a} and J = I_{b} are two intervals of ⋃_{n} ℱ_{n}, let IJ = I_{ab}. Suppose first that there exists κ > 0 such that

(18)  \frac{1}{κ}σ(I) ≤ ω(I) ≤ κσ(I)  for all I ∈ ⋃_{n≥0} ℱ_{n}.

Since every open set of [0, 1) is a countable disjoint union of intervals of ⋃_{n} ℱ_{n}, this inequality can be extended to any open set. By regularity of the measures σ and ω, it is also true for every Borel set. Thus, (i) is satisfied.

If (18) is not satisfied, we can for example suppose that there exists an integer n_{0} ≥ 0 and an interval I_{0} ∈ ℱ_{n_{0}} such that

ω(I_{0}) < \frac{1}{lC} σ(I_{0})

where C is the constant which appears in (16). It follows that for every I ∈ ⋃_{n} ℱ_{n},

\frac{ω(II_{0})}{ω(I)} ≤ \frac{1}{l}σ(I_{0}) = l^{−(n_{0}+1)}.

The related mass of the subinterval II_{0} is then smaller than expected and we can use the same idea as in [H3] and [H5] to prove that the dimension of the measure ω is strictly less than 1. This idea was also previously used by Bourgain in [Bg] and Batakis in [Ba]. For every I ∈ ⋃_{n} ℱ_{n} and for every x ∈ [0, 1), we have

\begin{aligned}∑_{J∈ℱ_{n_{0}}} ω(IJ)^{x} & ≤ ω(II_{0})^{x} + (l^{n_{0}} − 1)[\frac{ω(I) − ω(II_{0})}{l^{n_{0}} − 1}]^{x} \\ & ≤ (l^{−(n_{0}+1)x} + (l^{n_{0}} − 1)[\frac{1 − l^{−(n_{0}+1)}}{l^{n_{0}} − 1} ]^{x})ω(I)^{x} = β(x)ω(I)^{x}. \end{aligned}

If we sum over all intervals I of the same generation and iterate this inequality, we obtain

∑_{I∈ℱ_{pn_{0}}} ω(I)^{x} ≤ (β(x))^{p}ω([0, 1))^{x} for all p ≥ 0.

Using definition (3), we conclude that

τ_{pn_{0}}(x) ≤ \frac{1}{n_{0} log l} log β(x)  and  τ(x) ≤ \frac{1}{n_{0} log l} log β(x).

As recalled in the introduction, the function τ is differentiable and the dimension of the measure is equal to −τ′(1). Observing that β(1) = 1 and using the strict convexity of the logarithm, it is easy to establish that

dim (ω) = −τ′(1) ≤ −\frac{1}{n_{0} log l} β′(1) < 1.

The case where ω(I_{0}) > lCσ(I_{0}) can be treated in a similar way.

The end of the proof of Corollary 3.2 is easy. The function τ satisfies

τ(0) = 1,  τ(1) = 0  and  − τ′(1) = d < 1.

Using the convexity of τ we deduce that

−τ′(0) = δ > d.

Let τ^{∗} be the Legendre transform of τ defined by

τ^{∗}(α) = inf{αx + τ(x) : x ∈ R}.

It is well known that the function τ^{∗} is increasing in (−τ′(+∞), −τ′(0)] and then decreasing in [−τ′(0), −τ′(−∞)). On the other hand, if I_{n}(t) is the unique interval I ∈ ℱ_{n} such that t ∈ I and if E_{α} is defined as in (17), the doubling property ensures that

E_{α} = \{t ∈ [0, 1) : lim_{n→+∞} \frac{ω(I_{n}(t))}{−n log l} = α\}.

Using [BMP] and the differentiability of the function τ, we get

dim (E_{α}) = τ^{∗}(α) ≥ τ^{∗}(d) = d for all α ∈ [d, δ].

The doubling property clearly implies that d > 0 (in fact we will see in Section 5 that d ≥ \frac{1}{2}(1 + ε)). Then, we can conclude that dim (E_{α}) is positive for infinitely many values of the parameter α.

4. Characterization of the functions g such that dim (ω) < 1

In this section, we shall describe the set of Lipschitz functions g such that dim (ω) < 1. In particular, we prove that its complement is closed and nowhere dense in the set of Lipschitz functions. With the same notation as in Section 3, we obtain the two following results.

Theorem 4.1. There are only two possible cases, mutually exclusive.

(i) f is Lipschitz with ∥f′∥_{∞} ≤ ∥g′∥_{∞}/(\sqrt{l} − 1). In this case, ω is strongly equivalent to the Lebesgue measure and dim (ω) = 1;

or (ii) there exist two positive numbers a and b such that for every interval I of

length |I| ≤ 1,

(19)  a|I|^{1/2} ≤ osc (f, I) ≤ b|I|^{1/2},

where osc (f, I) = max_{I} f − min_{I} f. In this case, ω is singular with respect to the Lebesgue measure and dim (ω) < 1.

Theorem 4.2. The set of Lipschitz functions g such that f is Lipschitz is a closed vector subspace of infinite codimension of the space of 1-periodic Lipschitz functions (endowed with its natural topology). Consequently, its complement is an open dense subset of the space of 1-periodic Lipschitz functions.

Remarks. 1. Theorems 4.1 and 4.2 ensure that the case where dim (ω) < 1 is generic.

2. Conclusion (19) is often present in the literature (note that the right-hand part of (19) simply means that f is of class C^{1/2} ). Falconer states that (19) is sufficient to prove that the box-counting dimension of the graph ∂Ω is equal to \frac{3}{2} (see [F]). In [PU], the authors find sufficient conditions on g which ensure that (19) is satisfied. Finally, assuming some stronger regularity properties on g, Kaplan et al. [KMPY] prove a theorem similar to ours about the behaviour of f.

This section is organized as follows. In the first subsection, we obtain a geometric condition (on the function f) which ensures that dim (ω) < 1. In Section 4.2, we describe the behaviour of the function f and prove Theorem 4.1. In Section 4.3, we establish a necessary and sufficient condition for the functional equation f(t) = g(t) + f(lt)/\sqrt{l} to have a Lipschitz solution and then prove Theorem 4.2. Finally, in the last subsection, we give an example where f is not Lipschitz and therefore dim (ω) < 1.

4.1. A geometric condition on f which implies that dim (ω) < 1

Proposition 4.3. Suppose that f is C^{1/2} (not necessarily of Weierstrasstype). If we can find t_{0} ∈ [0, 1) and r > 0 such that

(20) f(t_{0} + h) ≥ f(t_{0}) for all h ∈ [0, r) and lim sup_{h→0^{+}} \frac{f(t_{0} + h) − f(t_{0})}{h^{1/2}} > 0,

then

lim_{s→0} \frac{ω^{M_{∅}}(Δ_{s})}{s} = 0,

where Q_{0} = (f(t_{0}), t_{0}) and Δ_{s} = Δ(Q_{0}, s).

Proposition 4.3 is not specific to Weierstrass-type curves; it is true for any function of class C^{1/2} , but does not imply dim (ω) < 1 in general. However, in the case where f is a Weierstrass-type function, the existence of one point satisfying (20) implies the existence of many other such points where the behaviour of the measure is similar. The rigorous property behind this heuristic argument is the quasi-Bernoulli property for the measure ω and, as a consequence of Corollary 3.2, we have

Corollary 4.4. Suppose that f is a Weierstrass-type function. Under condition (20), we have

dim (ω) < 1.

Remark. The conclusion of Corollary 4.4 would also hold if we could find t_{0} ∈ [0, 1) and r > 0 such that

(21) f(t_{0} + h) ≤ f(t_{0}) for all h ∈ [0, r) and lim inf_{h→0^{+}} \frac{f(t_{0} + h) − f(t_{0})}{h^{1/2}} < 0.

In that case, we could prove that

lim_{s→0} \frac{ω^{M_{∅}}(Δ_{s})}{s} = +∞.

Proof of Proposition 4.3. Let δ_{s} be the projection on the graph of [t_{0}, t_{0} + s). We know that ω^{M_{∅}}(Δ_{s}) ≈ ω^{M_{∅}}(δ_{s}). The maximum principle ensures that ω^{M_{∅}}(δ_{s}) does not depend on the values of f(t) for t < t_{0} . So, we can suppose that f(t) = f(t_{0}) for every t < t_{0} . On the other hand, if M̃_{∅} = Q_{0} + (10Kr^{1/2}, r), we have previously remarked that ω^{M_{∅}}(Δ_{s}) ≈ ω^{_{M}^{∅}}(Δ_{s}) when s → 0 (here, the constants may depend on r). Since ω^{_{M}^{∅}}(Δ_{s}) does not depend on the values of f(t) for t ≥ t_{0} + r, we can also suppose that f(t) = f(t_{0} + r) for every t ≥ t_{0} + r. In other words, these two remarks say that we can suppose that t_{0} is a global minimum for the function f.

Let Ω̂ = {(x, t) : x > f(t_{0})}. We have Ω ⊂ Ω̂. Therefore, if G and Ĝ denote the Green functions on Ω and Ω̂, we know that

(22)  G(M, P) = Ĝ(M, P) − R̂_{_{G}(M,·)}^{_{Ω}\Ω}(P),

where R̂_{_{G}(M,·)}^{_{Ω}\Ω} is the smooth reduction of Ĝ(M, ·) on Ω̂ \ Ω (for more details about reduction, we can refer to [Do]). Let ω̂^{M_{∅}} be the caloric measure at M_{∅} related to Ω̂ and Δ̂_{s} = {(f(t_{0}), t) : t ∈ (t_{0} − s, t_{0} + s)}. Using Theorem 1.5 we get

\frac{ω^{M_{∅}}(Δ_{s})}{s} ≈ \frac{ω^{M_{∅}}(Δ_{s})}{ω̂^{M_{∅}}(Δ̂_{s})} ≈ \frac{G(Q_{s}, M_{∅})}{Ĝ(Q_{s}, M_{∅})}

where Q_{s} = Q_{0} + (\sqrt{s}, 0).

To complete the proof of the proposition, we adapt to our situation an old idea due to Naïm ([Na]). We must be careful, because the heat operator is not selfadjoint. According to [K] and [H2], let χ be the unique caloric minimal function in Ω̂ which tends to 0 at every point of ∂Ω̂ \ {Q_{0}} and satisfies χ(M_{∅}) = 1. Following [H2, Part 4], we know that

χ(P) = lim_{s→0} \frac{Ĝ(Q_{s}, P)}{Ĝ(Q_{s}, M_{∅})}.

On the other hand, the additivity and positivity of the map s ↦ R̂_{s}^{_{Ω}\Ω}(M_{∅}) ensure the existence of a non-negative Radon measure μ such that for every non-negative super-caloric function,

R̂_{s}^{_{Ω}\Ω}(M_{∅}) = ∫ s(ξ) dμ(ξ).

Using (22) and Fatou’s lemma, we get

lim sup_{s→0} \frac{G(Q_{s}, M_{∅})}{Ĝ(Q_{s}, M_{∅})} = 1 − lim inf_{s→0} ∫ \frac{Ĝ(Q_{s}, ξ)}{Ĝ(Q_{s}, M_{∅})} dμ(ξ)

≤ 1 − ∫ χ(ξ) dμ(ξ) = 1 − R̂_{χ}^{_{Ω}\Ω}(M_{∅}).

According to [H2, Proposition 4.2], hypothesis (20) implies that the set Ω̂ \ Ω is not thin in the minimal function χ. So,

R̂_{χ}^{_{Ω}\Ω}(M_{∅}) = χ(M_{∅}) = 1.

This completes the proof.

4.2. Proof of Theorem 4.1. In this section, we identify 1-periodic functions with functions defined on the torus R/Z. The usual distance on R/Z is noted d. We begin with the following elementary lemma.

Lemma 4.5. Suppose that ∥g′∥_{∞} ≤ 1. Let s, t ∈ R/Z such that d(t, s) ≤ 1/2l (this is equivalent to d(lt, ls) = l d(t, s)) and suppose

\frac{|f(lt) − f(ls)|}{d(lt, ls)} ≥ \frac{1}{\sqrt{l} − 1}(1 + u) for some u ≥ 0.

Then

\frac{|f(t) − f(s)|}{d(t, s)} ≥ \frac{1}{\sqrt{l} − 1}(1 + u\sqrt{l}).

The above lemma is an easy consequence of the functional equation between f and g. We have

f(t) − f(s) = g(t) − g(s) + \frac{1}{\sqrt{l}} (f(lt) − f(ls)).

Thus,

\frac{|f(t) − f(s)|}{d(t, s)} ≥ \sqrt{l} \frac{|f(lt) − f(ls)|}{d(lt, ls)} − \frac{|g(t) − g(s)|}{d(t, s)} ≥ \sqrt{l} \frac{|f(lt) − f(ls)|}{d(lt, ls)} − 1

which can be rewritten as

\frac{|f(t) − f(s)|}{d(t, s)} − \frac{1}{\sqrt{l} − 1} ≥ \sqrt{l} [\frac{|f(lt) − f(ls)|}{d(lt, ls)} − \frac{1}{\sqrt{l} − 1}]

and the lemma is proved.

Iterating Lemma 4.5, we obtain the following corollary.

Corollary 4.6. Suppose once again that ∥g′∥_{∞} ≤ 1. Let s, t ∈ R/Z such that d(t, s) ≤ 1/2l^{n} (this is equivalent to d(l^{n}t, l^{n}s) = l^{n} d(t, s)) and suppose

\frac{|f(l^{n}t) − f(l^{n}s)|}{d(l^{n}t, l^{n}s)} ≥ \frac{1}{\sqrt{l} − 1}(1 + u) for some u ≥ 0.

Then

\frac{|f(t) − f(s)|}{d(t, s)} ≥ \frac{1}{\sqrt{l} − 1}(1 + ul^{n/2}).

Now we can prove Theorem 4.1. Let us assume, without loss of generality, that the Lipschitz constant of g is equal to 1. If f is Lipschitz, it is known that the measure ω is strongly equivalent to the Lebesgue measure (see [H2, pp. 642–643]).

On the other hand, suppose that f is not (\sqrt{l} − 1)^{−1} -Lipschitz. This means that we can find two points s_{0}, t_{0} ∈ R/Z such that

\{ \begin{matrix}\frac{|f(t_{0}) − f(s_{0})|}{d(t_{0}, s_{0})} = \frac{1}{\sqrt{l} − 1}(1 + v), & v > 0, \\ t_{0} = s_{0} + α, & 0 < α ≤ \frac{1}{2}. \end{matrix}

Let I be a closed interval on the circle R/Z and let n be the unique integer ≥ 1 such that

l^{−n} ≤ \frac{1}{2}|I| < l^{−n+1}.

We can choose s_{n} such that s_{n} ∈ I, s_{n} + l^{−n} ∈ I and l^{n}s_{n} = s_{0} . Define t_{n} = s_{n} + l^{−n}α; we have l^{n}t_{n} = t_{0} and t_{n} ∈ I. Using Corollary 4.6 we obtain

\frac{|f(t_{n}) − f(s_{n})|}{d(t_{n}, s_{n})} ≥ \frac{1}{\sqrt{l} − 1}(1 + vl^{n/2}) ≥ \frac{vl^{n/2}}{\sqrt{l} − 1}.

We conclude that

osc (f, I) ≥ \frac{v d(t_{0}, s_{0})}{\sqrt{l} − 1}l^{−n/2} ≥ a|I|^{1/2}

with a = v d(t_{0}, s_{0})/(\sqrt{l} − 1)(2l)^{1/2} .

This inequality ensures that (20) is satisfied if t_{0} is the minimum of f. Then, Corollary 4.4 allows us to conclude that dim (ω) < 1.

4.3. Proof of Theorem 4.2. We know that, when g is Lipschitz, the following functional equation

(23)  f(t) = g(t) + \frac{1}{\sqrt{l}} f(lt)  for all t ∈ R/Z,

(qua equation in f) has a unique solution in C^{0}(R/Z, R), defined by

f(t) = ∑_{n=0}^{∞} l^{−n/2}g(l^{n}t).

We would like to know if f is Lipschitz or, equivalently, if (23) has a solution in Lip (R/Z, R).

Proposition 4.7. Let g ∈ Lip (R/Z, R) and f ∈ C^{0}(R/Z, R) satisfying (23). Define the linear operator A_{l} from Lip (R/Z, R) into itself by

A_{l}(φ) = ψ with ψ(t) = \frac{1}{l} ∑_{ls=t} φ(s)

and put U_{l} = −(Id −\sqrt{l} A_{l})^{−1}(\sqrt{l} A_{l}) (the operator Id −\sqrt{l} A_{l} is invertible). Then, f is Lipschitz if and only if f = U_{l}(g). Equivalently, f is Lipschitz if and only if

(24)  U_{l}(g)(t) = g(t) + \frac{1}{\sqrt{l}} U_{l}(g)(lt)  for all t ∈ R/Z.

This is a closed, linear condition, of infinite codimension in Lip (R/Z, R).

Proof. Note that equation (23) implies the following (a priori weaker) condition

(25)  A_{l}(f − g) = \frac{1}{\sqrt{l}} f.

We claim that (25), qua equation in f, has a unique solution in Lip (R/Z, R). Therefore, (23) will have a solution in Lip (R/Z, R) if and only if the Lipschitz solution of (25) also verifies (23).

Equation (25) can be rewritten as

(Id −\sqrt{l} A_{l})f = −\sqrt{l} A_{l}g.

To prove our claim, we have to establish that the operator Id −\sqrt{l} A_{l} is invertible on Lip (R/Z, R). This will imply that (25) has a unique solution, namely

U_{l}(g) = −(Id −\sqrt{l} A_{l})^{−1}(\sqrt{l} A_{l})(g)

and will give the first part of Proposition 4.7.

To prove that Id −\sqrt{l} A_{l} is invertible, let us denote by 1 the constant function equal to 1 and let

Lip^{0}(R/Z, R) = {g ∈ Lip (R/Z, R) : ∫_{R/Z} g(t) dt = 0}.

We have the decomposition

Lip (R/Z, R) = Lip^{0}(R/Z, R) ⊕ R1

which is preserved by \sqrt{l} A_{l}. The operator Id −\sqrt{l} A_{l} is obviously invertible on R1, so we simply have to prove invertibility on Lip^{0}(R/Z, R). But, \sqrt{l} A_{l} is a contraction on Lip^{0}(R/Z, R) (endowed with norm ∥g′∥_{∞}). So, Id −\sqrt{l} A_{l} is invertible in this space and its inverse is given by ∑_{k≥0}(\sqrt{l} A_{l})^{k}. Moreover, we have ∥(Id −\sqrt{l} A_{l})^{−1}∥ ≤ 1/(1 − l^{−1/2}) on Lip^{0}(R/Z, R). Therefore, Id −\sqrt{l} A_{l} is indeed invertible on the whole space Lip (R/Z, R). This proves the claim.

To finish the proof of Proposition 4.7, we observe that condition (24), which states that U_{l}(g) verifies (23), is closed and linear. Moreover, it is easy to construct a lot of functions g which do not verify (24). An elementary calculation gives

U_{l}(1) = \frac{1}{1 − l^{−1/2}} and U_{l}(e^{2iπnt}) = − ∑_{k≥1, l^{k}|n} l^{k/2}e^{2iπ(n/l^{k})t} if n ≠ 0.

In particular, when g is a nonzero trigonometric polynomial whose frequencies are in Z \ lZ, condition (24) is not verified and the Weierstrass-like function f is not Lipschitz.

More generally, when g is Lipschitz, we can compute the Fourier coefficients of U_{l}(g) and we obtain that g satisfies (24) if and only if, for every n ∈ Z,

(26)  l |̸ n ⟹ ∑_{k=0}^{+∞} l^{k/2}ĝ(l^{k}n) = 0.

These formulas have been already established by Kaplan et al. ([KMPY]) under more restrictive hypotheses on g.

All these remarks allow us to construct an infinite-dimensional space of functions g which do not verify condition (24), except the null function. In other words, the space of functions g such that f is Lipschitz has infinite codimension.

To conclude, note that the estimate

∥(Id −\sqrt{l} A_{l})^{−1}∥ ≤ \frac{1}{1 − l^{−1/2}} on Lip^{0}(R/Z, R)

implies that for every g ∈ Lip (R/Z, R),

∥U_{l}(g)′∥_{∞} ≤ \frac{1}{\sqrt{l} − 1}∥g′∥_{∞}.

If f is Lipschitz (that is if f = U_{l}(g)), this gives another proof of the inequality

∥f′∥_{∞} ≤ \frac{1}{\sqrt{l} − 1}∥g′∥_{∞}.

4.4. An example. To finish this section, we give a simple sufficient condition on g and l which guarantees dim (ω) < 1.

Proposition 4.8. Suppose that g is not constant. If l is sufficiently large, then

dim (ω) < 1.

In fact, it happens as soon as

\sqrt{l} \frac{\sqrt{l} − 1}{\sqrt{l} + 1} > \frac{∥g′∥_{∞}}{2osc (g, R)}.

Proof. Assume f and g are both Lipschitz. Then,

(27)  \{ \begin{matrix}f(t) = g(t) + f(lt)/\sqrt{l}, \\ f′(t) = g′(t) + \sqrt{l}f′(lt) & almost everywhere. \end{matrix}

We can easily deduce that

(28)  \{ \begin{matrix}osc (g, R) ≤ [1 + \frac{1}{\sqrt{l}} ]osc (f, R), \\ ∥f′∥_{∞} ≤ \frac{∥g′∥_{∞}}{\sqrt{l} − 1} \end{matrix}

(the second inequality is not new; we have already proved it, by two different methods in the previous sections).

Observing that osc (f, R) ≤ \frac{1}{2}∥f′∥_{∞} (since f is 1-periodic), we get

osc (g, R) ≤ \frac{\sqrt{l} + 1}{2\sqrt{l} (\sqrt{l} − 1)}∥g′∥_{∞}

and the proof of Proposition 4.8 is complete.

5. A minoration for the lower dimension of caloric measure

In this section, f is a bounded function of class C^{1/2} (not necessarily of Weierstrass-type) with Hölder constant K. Denote by Ω the corresponding domain and choose an arbitrary point M_{0} in Ω. We want to give estimates for the lower dimension of ω̃^{M_{0}} , which is defined by

dim_{∗}(ω̃^{M_{0}}) = inf(dim (E) : ω̃^{M_{0}}(E) > 0).

Using Taylor–Watson results about the description of polar sets related to the heat equation ([TW]), we know that dim_{∗}(ω̃^{M_{0}}) is not less than \frac{1}{2} . In the next theorem, we improve this result.

Theorem 5.1. There exists ε = ε(K) > 0 such that for every M_{0} ∈ Ω,

dim_{∗}(ω̃^{M_{0}}) ≥ \frac{1}{2}(1 + ε).

Remark. If dim^{∗}(ω̃^{M_{0}}) = inf(dim (E) : ω̃^{M_{0}}(R \ E) = 0), we obviously conclude that

dim^{∗}(ω̃^{M_{0}}) ≥ dim_{∗}(ω̃^{M_{0}}) ≥ \frac{1}{2}(1 + ε).

In fact, when f is a Weierstrass-type function, the quasi-Bernoulli property, proved in Section 3, ensures that

dim^{∗}(ω̃^{M_{0}}) = dim_{∗}(ω̃^{M_{0}}) = dim(ω̃^{M_{0}}).

However, for a general function f, we can have

dim^{∗}(ω̃^{M_{0}}) > dim_{∗}(ω̃^{M_{0}}).

Proof of Theorem 5.1. We first remark that if M_{0} = (x_{0}, t_{0}) and M_{1} = (x_{1}, t_{1}) are two points in Ω with t_{0} ≤ t_{1} , then, the measures ω^{M_{0}} and ω^{M_{1}} have the same null sets in ∂Ω ∩ {t ≤ t_{0}}. In fact, Harnack inequalities ensure that ω^{M_{0}} is absolutely continous with respect to ω^{M_{1}} and Theorem 1.1 ensures that the measure ω^{M_{1}} is absolutely continous with respect to ω^{M_{0}} in ∂Ω ∩ {t ≤ t_{0}}. To prove Theorem 5.1, it is then sufficient to estimate the Hausdorff dimension of non negligible sets E related to ω̃^{M_{0}} when E ⊂ Π(Δ(Q_{0}, \frac{1}{2}r)), M_{0} = Q_{0} + (20K\sqrt{r}, 8r) and r arbitrarily small (see (1) and (6) for the notation). We will use the following lemma.

Lemma 5.2. There are two positive numbers C = C(K) and ε = ε(K) such that for every Q_{0} = (f(t_{0}), t_{0}) ∈ ∂Ω and r > 0,

ω̃^{M_{0}}([s_{0} − s, s_{0} + s]) ≤ C(\frac{s}{r})^{(1+ε)/2},

where M_{0} = Q_{0} + (20K\sqrt{r}, 8r) and [s_{0} − s, s_{0} + s] ⊂ [t_{0} − r, t_{0} + r].

It is elementary to deduce from Lemma 5.2 that there exists C = C(r, K) > 0 such that for every Borel set E ⊂ [t_{0} − r, t_{0} + r],

ω̃^{M_{0}}(E) ≤ Cℋ ^{(1+ε)/2}(E)

where ℋ ^{(1+ε)/2} is the Hausdorff measure in dimension \frac{1}{2}(1 + ε). Then, the lower dimension of the measure ω̃^{M_{0}} restricted to [t_{0} −r, t_{0} +r] is not less than \frac{1}{2}(1+ε) and the theorem follows.

Proof of Lemma 5.2. Using (14) in the present situation, we have

ω̃^{M_{0}}([t − s, t + s]) ∼ < \sqrt{\frac{s}{r}} φ^{∗}(Q + (\sqrt{s}, 0))

where Q = (f(s_{0}), s_{0}) and φ^{∗} is the 𝒞^{∗} -harmonic measure of ∂T (Q, r) \ Δ(Q, r) in T (Q, r). Let us introduce

U = \{(x, t) ∈ R^{2} : |t| < 1, |x| < 10K and x > −10K\sqrt{|t|}\}

and denote by ψ^{∗} the 𝒞^{∗} harmonic measure of ∂U \ \{(x, t) : x = −10K\sqrt{|t|} \} in U. Using the maximum principle and the homogeneity of the operator 𝒞^{∗}, we have

φ^{∗}(Q + (\sqrt{s}, 0)) ≤ ψ^{∗}(\sqrt{\frac{s}{r}}, 0).

Therefore, Lemma 5.2 is a consequence of the following

Lemma 5.3. There are two positive numbers C = C(K) and ε = ε(K) such that

ψ^{∗}((x, 0)) ≤ Cx^{ε} for all x ∈ (0, 5K].

Proof. Similar ideas can be found in [H2]. Let us sketch the proof to be self-contained. Put

V = {(x, t) ∈ R^{2} : |t| < 1 and |x| < 10K}

and denote by G(A, ·) the Green function of V with singularity at A. Then, G( · , B) is the Green function of V related to the adjoint operator with singularity at B. Let B = (−8K, \frac{1}{2}) and observe that B ∈ V \ U. If

Ṽ = \{(x, t) ∈ R^{2} : |t| < \frac{1}{4} and |x| < 5K\},

there exists a constant C_{1} = C_{1}(K) > 0 such that

G(M, B) ≥ C_{1} for all M ∈ Ṽ.

Caloric measure on domains bounded by Weierstrass-type graphs 521

On the other hand, if Γ( · , ·) is the heat kernel, and if

W = B + {(x, t) ∈ R^{2} : |t| < 1/100 and |x| < K/10},

there exists a constant C_{2} = C_{2}(K) > 0 such that

(29)  G(M, B) ≤ Γ(M, B) ≤ C_{2}  for all M ∈ ∂W.

Using the maximum principle, we extend (29) to V \ W. In particular, it is true in U. Then, by the maximum principle,

1 − ψ^{∗}(M) ≥ \frac{1}{C_{2}} G(M, B) for all M ∈ U.

We can conclude that

ψ^{∗}(M) ≤ 1 − \frac{C_{1}}{C_{2}} for all M ∈ Ṽ ∩ U.

Iterating this inequality, we have

ψ^{∗}(M) ≤ (1 − \frac{C_{1}}{C_{2}} )^{k}

where k ≥ 1 and M ∈ U ∩ {(x, t) : |t| ≤ 4^{−k}, |x| ≤ 10K2^{−k}}. We can easily deduce that Lemma 5.3 is true with

ε = \frac{− log(1 − C_{1}/C_{2})}{log 2}.

References

[Ba] Batakis, A.: Harmonic measure of some Cantor type sets. - Ann. Acad. Sci. Fenn. Math. 21, 1996, 255–270.

[BH] Batakis, A., and Y. Heurteaux: On relations between entropy and Hausdorff dimension of measures. - Preprint, Orsay, 1998.

[Bg] Bourgain, J.: On the Hausdorff dimension of harmonic measure in higher dimension. - Invent. Math. 87, 1987, 477–483.

[Bo] Bousch, T.: Le poisson n’a pas d’arêtes. - Preprint, 1998.

[BMP] Brown, G., G. Michon, and J. Peyrière: On the multifractal analysis of measures. - J. Stat. Phys. 66, 1992, 775–790.

[C] Carleson, L.: On the support of harmonic measure for sets of Cantor type. - Ann. Acad. Sci. Fenn. Math. 10, 1985, 113–123.

[Da] Dahlberg, B.: Estimates of harmonic measure. - Arch. Rational Mech. Anal. 65, 1977, 275–288.

522 Thierry Bousch and Yanick Heurteaux

[Do] Doob, J.L.: Classical Potential Theory and its Probabilistic Counterpart. - Springer-Verlag, New York, 1984.

[FGS] Fabes, E.B., N. Garofalo, and S. Salsa: A backward Harnack inequality and Fatou theorem for nonnegative solutions of parabolic equations. - Illinois J. Math. 30, 1986, 536–565.

[F] Falconer, K.: Fractal Geometry. - Mathematical Foundations and Applications. John Wiley & Sons, New York, 1990.

[H1] Heurteaux, Y.: Inégalités de Harnack à la frontière pour des opérateurs paraboliques. - Ph.D. thesis, Université Paris-Sud, 1989.

[H2] Heurteaux, Y.: Solutions positives et mesure harmonique pour des opérateurs paraboliques dans des ouverts “lipschitziens”. - Ann. Inst. Fourier (Grenoble) 41, 1991, 601–649.

[H3] Heurteaux, Y.: Sur la comparaison des mesures avec les mesures de Hausdorff. - C. R. Acad. Sci. Paris Sér. I Math. 321, 1995, 61–65.

[H4] Heurteaux, Y.: Mesure harmonique et équation de la chaleur. - Ark. Mat. 34, 1996, 119–139.

[H5] Heurteaux, Y.: Estimations de la dimension inférieure et de la dimension supérieure des mesures. - Ann. Inst. H. Poincaré Probab. Statist. 34, 1998, 309–338.

[KMPY] Kaplan, J.L., J. Mallet-Paret, and J.A. Yorke: The Lyapunov dimension of a nowhere differentiable attracting torus. - Ergodic Theory Dynam. Systems 4, 1984, 261–281.

[KW] Kaufman, R., and J.M. Wu: Parabolic measure on domains of class Lip 1/2. - Compositio Math. 65, 1988, 201–207.

[K] Kemper, J.T.: Temperatures in several variables: kernel functions, representations and parabolic boundary values. - Trans. Amer. Math. Soc. 167, 1972, 243–262.

[LS] Lewis, J.L., and J. Silver: Parabolic measure and the Dirichlet problem for the heat equation in two dimensions. - Indiana Univ. Math. J. 37, 1988, 801–839.

[MV] Makarov, N., and A. Volberg: On the harmonic measure of discontinuous fractals. - Preprint LOMI E-6-86, Leningrad, 1986.

[Na] Naïm, L.: Sur le rôle de la frontière de R.S. Martin dans la théorie du potentiel. - Ann. Inst. Fourier (Grenoble) 7, 1957, 183–281.

[Ny] Nyström, K.: The Dirichlet problem for second order parabolic operators. - Indiana Univ. Math. J. 46, 1997, 183–245.

[PU] Przytycki, F., and M. Urbanski: On the Hausdorff dimension of some fractal sets. - Studia Math. 93, 1989, 155–186.

[TW] Taylor, S.J., and N.A. Watson: A Hausdorff measure classification of polar sets for the heat equation. - Math. Proc. Cambridge Philos. Soc. 97, 1985, 325–344.

[W] Wu, J.M.: On parabolic measures and subparabolic functions. - Trans. Amer. Math. Soc. 251, 1979, 171–186.

Received 14 June 1999